### v1

# Water Detection in the Interstellar Object 3I/ATLAS

Abstract
We report the first detection of water activity in the third confirmed interstellar object, 3I/ATLAS, based on ultraviolet imaging with the Neil Gehrels-Swift Observatory. Observations acquired with the Ultraviolet/Optical Telescope on 2025 July 31st - Aug 1st revealed OH (A — X) emission near 3085 Å. The water production rate results highly depend on the reddening assumption. For a reddening of 38.6% between 5437.8 Å and 3325.7 Å, the water production rate is molecules s-1 (40 kg s-1) at a heliocentric distance of 3.51 au. This places 3I/ATLAS among the few comets with confirmed OH emission beyond 3 au, where water ice sublimation is typically inefficient. The inferred production rate is consistent with an active area of at least 19 km2, assuming equilibrium sublimation. Based on current upper limits of the nucleus’ radius, this requires that over 20% of the surface is active, which is larger than activity levels observed in most solar system comets. Contemporaneous near-infrared spectroscopy indicates the presence of large icy grains in the coma, which may serve as an extended source of water vapor. The detection of OH emission prior to any CN detection is unusual and may reflect differences in grain-driven outgassing or volatile inventory compared to typical comets. While similar behavior has been observed in solar system comets, the mechanisms controlling distant activity and the storage and release of volatiles remain poorly understood. If 3I/ATLAS’ coma continues to be dominated by H2O, supporting the early and low-metallicity formation hypothesis, the derived large size of the nucleus could be indicative of a key knowledge gap in low-metallicity system planetesimal formation and loss mechanisms.
1 Introduction
The discovery of the third interstellar object, 3I/ATLAS, on 1 July 2025 initiated a broad characterization campaign across the globe (L. Denneau, 2025; D. Z. Seligman et al., 2025). Following the lessons learned from the prior interstellar objects 1I/‘Oumuamua and 2I/Borisov observing campaigns were initiated to rapidly capture its initial brightness, morphology, lightcurves, color, and optical and near-infrared spectrum (‘Oumuamua Team, 2019; A. Fitzsimmons et al., 2023). Given the apparent brightness and early extension of the coma the production of gas was assumed and searched for (A. Alvarez-Candal et al., 2025; C. Opitom et al., 2025) but not found. Characterizing the early activity of interstellar objects is essential for understanding their chemical and physical evolution during solar approach, as it possibly represents the first time the object has been significantly heated in their very long dynamic lifetimes (M. J. Hopkins et al., 2025; A. G. Taylor & D. Z. Seligman, 2025).
Inbound activity for our own Oort Cloud solar system comets is often more predictable. Hypervolatiles like CO and CO2 drive activity beyond the water ice line (T. Ootsubo et al., 2012; O. H. Pinto et al., 2022), which can entrain small icy grains into the coma that sublimate to produce H2O (M. F. A’Hearn et al., 2011). As comets approach the Sun, water increasingly dominates the observed gas composition. This is not necessarily true for all solar system comets (see C/1908 R1 (Morehouse), C/2016 R2 (PANSTARRS), and C/2009 P1 (Garradd); (A. J. McKay et al., 2019; N. Biver et al., 2018; D. Bodewits et al., 2014), nor was it the case for interstellar comet 2I/Borisov, which transitioned from a equally mixed H2O/CO coma pre-perihelion to a CO-dominated coma post-perihelion (D. Bodewits et al., 2020; Z. Xing et al., 2020; M. A. Cordiner et al., 2020). The change in composition was only captured thanks to the fortunate timing of UV observations by HST and the Neil Gehrels-Swift Observatory that monitored CO and H2O production. A similar monitoring campaign was planned for the next interstellar object to establish a chemical baseline for comparison.
The Ultraviolet and Optical Telescope (UVOT) onboard the Neil Gehrels-Swift Observatory has proven to be an excellent instrument for monitoring the OH (A – X) band at 3085 Å (K. Mason et al., 2007; J. A. Carter et al., 2012; D. Bodewits et al., 2014; Z. Xing et al., 2020; D. Bodewits et al., 2023). Despite its modest aperture, UVOT’s sensitivity is enhanced by its position above Earth’s atmosphere, free from telluric absorption and sky background, enabling deep integrations and effective stacking. Because it is flown above the atmospheric extinction, UVOT has the equivalent sensitivity of a 4-m telescope on the ground111https://swift.gsfc.nasa.gov/proposals/tech_appd/swiftta_v12.pdf. UVOT was used to derive the H2O production rate of 2I/Borisov during its flyby of the solar system in 2019/2020, and it was the dual measurements of UVOT and HST COS that showed the duality of Borisov’s chemical composition around perihelion (D. Bodewits et al., 2020; Z. Xing et al., 2020).
Here, we present our initial UVOT observations of 3I/ATLAS taken on July 31 and August 1, 2025 that show the onset of water activity in the third interstellar object. In the following sections we describe the observations and analysis, place our results in context with the current 3I/ATLAS literature, and compare and contrast with both the previous interstellar objects and Oort Cloud comet populations.
2 Observation and Analysis
We observed 3I/ATLAS with the Ultraviolet/Optical Telescope (UVOT; P. W. Roming et al., 2000) on board the Neil Gehrels-Swift Observatory (N. Gehrels et al., 2004). UVOT is a 30-cm Ritchey–Chrétien telescope with a 17 arcmin17 arcmin field of view and a plate scale of 0.502 arcsec/pixel, although most images are taken in ‘image mode’ with 22 on-board binning, resulting in an effective plate scale of 1.004 arcsec/pixel. It is equipped with 11 broadband filters and two grisms, covering wavelengths from 1600 to 8000 Å.
Observations were obtained between 2025 July 31 05:24 UTC and 2025 August 01 09:49 UTC, when the comet was at a heliocentric distance of 3.51 au and a geocentric distance of 2.82 au (Table 1). We used UVOT’s UVW1 filter (central wavelength = 2600 Å, FWHM = 693 Å), which includes most of the OH (A - X) emission band centered near 3085 Å. To estimate and subtract the continuum contribution within the UVW1 bandpass, we also acquired images using the V-band filter ( = 5468 Å, FWHM = 769 Å).
Swift does not track moving objects, so comets are typically observed with brief 200-s exposures. During our observations, 3I had a relatively high apparent motion of 1.79 to 1.86 arcsec/min, corresponding to a motion of about 6 arcsec in a typical 200-s exposure and thus exceeding UVOT’s point spread function (PSF; 2.18 arcsec FWHM in V and 2.37 arcsec FWHM in UVW1 according to A. A. Breeveld et al. (2010)). To avoid blurring and further dilution of the surface brightness, we acquired the UVW1 images in ‘event-mode’, where the arrival time of every photon is time tagged. Due to the telemetry load constraints of Swift and the high brightness in the V band images, we could not use event mode for the V band observations and instead acquired short-exposure images in the ‘image mode’.
| Mid Time (UTC) | (AU) | (km s-1) | (AU) | S–T–O (deg) | (days) | Filter | Images | Exp. Time (s) | ||||
[…]
To mitigate motion blur, we divided photons in each event-mode UVW1 image into 30-second time slices, limiting motion within each slice to less than 1 arcsec. We then aligned all slices by the target’s nucleus position (from JPL/Horizons) and summed them to produce motion-corrected UVW1 images. These UVW1 images have a pixel scale of 0.502 arcsec pixel-1, and we binned them by 2×2 to match the 1.004 arcsec pixel-1 scale of V-band images obtained in the image mode for subsequent image subtraction. For V-band images, such motion correction was not possible without event-mode images.
Next, we discarded all images in which the target was contaminated by background stars, then aligned the remaining images on the nucleus’ position and stacked them separately for each filter to improve the SNR ratio. The resulting UVW1 and V-band stacked images, shown in Fig. 1, have equivalent exposure times of 9055 s and 2713 s, respectively. Both images show clear detections of 3I/ATLAS. In the V-band image, the comet appears elongated along the direction of motion, as expected. The UVW1 image, free from smearing, shows a spatially extended source with a FWHM of 5 arcseconds and a slight asymmetry across the comet–Sun axis.
In the stacked images, we noticed that the target positions deviated from JPL/Horizons predictions. This offset, also present for stars in raw images, likely stems from a small error in the astrometric solution produced by the UVOT pipeline. The UVW1 stacked image showed particularly large errors (approximately 5 arcsec), suggesting greater uncertainty in event mode. We fitted the target with a 2D Gaussian to locate the nucleus position for further analyses.
To identify stars around the target, we applied a 3-sigma detection threshold requiring at least 4 connected pixels. The background was measured using a star-masked annulus with an inner radius of 30 arcsec and an outer radius of 40 arcsec (corresponding to 60,000 - 80,000 km). The derived backgrounds are counts s-1 pixel-1 for the V-band and counts s-1 pixel-1 for UVW1. After background subtraction, aperture photometry with a 10-arcsec radius yielded counts s-1 for UVW1, while an elongated 10-arcsec aperture accounting for motion blur yielded counts s-1 for V-band.
We converted the V-band net count rate to flux and magnitude using Swift’s effective area (corrected for UVOT’s ongoing sensitivity decline; A. A. Breeveld et al. (2011)) and a target spectrum model. We adopted a linearly reddened solar spectrum to model the target spectrum with the default solar spectrum included in the sbpy package (M. Mommert et al., 2019), and used a radius of 15 cm to derive the UVOT collecting area. The calculated flux, together with the same target spectrum model and Vega spectrum (R. C. Bohlin, 2014), was used to derive the apparent Vega magnitude in the V band of UVOT. For the V filter, the derived V-band fluxes and magnitudes are nearly insensitive to different reddening assumptions, with result variations less than 0.4%. Therefore, the uncertainty introduced by the reddening assumption is negligible for the photometry. The measured V-band flux is erg s-1 cm-2, corresponding to a V magnitude of mag.
The net count rate of the V image also enables an estimate of the reflected continuum contribution to the UVW1 flux, allowing us to isolate the OH (A – X) emission:
where is the count rate attributable to OH, and are the count rates measured in the UVW1 and V images, respectively, as described above, and is the solar continuum count rate ratio between the filters. The derived OH count rate depends heavily on the ratio factor , which is a function of the reddening per 1000 Å between the photon-weighted effective wavelengths of V and UVW1 filters. These effective wavelengths, calculated using the sensitivity-corrected effective area and a solar spectrum, are 5437.8 Å for V and 3325.7 Å for UVW1. To determine the count rate ratio , we convolved a solar spectrum linearly reddened by with the respective effective areas.
The OH count rates are converted to the emitted flux using a hydroxyl spectral model (D. Bodewits et al., 2019), then to luminosity using the geocentric distance. We subsequently derived the total number of OH molecules within the 10-arcsec radius aperture using fluorescence efficiencies (D. G. Schleicher & M. F. A’Hearn, 1988) appropriate for the specific geocentric velocities, with a scaling applied for the heliocentric distance.
To derive water production rates from the total number of OH molecules in the aperture, we applied the vectorial model with a random OH ejection kick of 1.05 km s-1 by taking photodissociation of H2O into account (M. C. Festou, 1981). We adopted a H2O outflow velocity of km s-1, photodissociation lifetimes of 86 000 s for H2O and 129 000 s for OH, and a branching ratio of 0.93 for water dissociation into OH (M. R. Combi et al., 2004). We compared observed and model-predicted OH molecular numbers within the same aperture. The observed water production rate was then obtained by scaling the modeled water production rate with this ratio.
3 Results and Discussion
The UVOT-V magnitude of mag corresponds to a Johnson V-mag of . This results in an of 174 cm after correcting for the phase angle (M. F. Ahearn et al., 1984; D. G. Schleicher & A. N. Bair, 2011). If we apply the V-R color of 0.51 mag measured on July 23 from T. Santana-Ros et al. (2025), the is 180 cm in the R band. The discrepancy between our results and the 300 cm reported by T. Santana-Ros et al. (2025) could originate in temporal coma development or aperture size differences.
Our water production rate measurement depends strongly on the assumed reddening between the effective wavelengths of the V and UVW1 filters (5437.8 Å and 3325.7 Å, respectively), as shown in the left panel of Fig. 2. The error bars represent 1- stochastic uncertainties. A possible water detection emerges for reddening assumptions larger than , while a clear 3- detection is achieved for reddening larger than .
We then investigated reddening measurements of 3I/ATLAS from other telescopes, presented in the middle panel of Fig. 2. The x-axis shows reddening values (percent per 1000 Å), while the y-axis indicates the central wavelength of the measurement bands. Different marker shapes denote literature sources, with colors representing the comet’s heliocentric distance during the observations (colorbar on the far right). A. Alvarez-Candal et al. (2025) measured per 1000 Å across 3200–5600 Å, with a central wavelength closest to ours (4381.8 Å). We show this reddening measurement in the left panel with a green dashed line. Other reddening measurements typically range from 10% to 25%, with higher reddening values at shorter wavelengths.
This trend of increasing reddening toward shorter wavelengths is unsurprising. For example, a trend of increased reddening at shorter central wavelengths is presented in D. Jewitt & K. J. Meech (1986). Directly measured values in the UV band corroborate this trend, with values of 60–80%/1000 Å near 2950 Å from C/1983 O1 (Černis) (P. D. Feldman & M. F. A’Hearn, 1985) and 70%/1000 Å near 2950 Å from 1P/Halley (P. D. Feldman et al., 1987). In contrast, ground-based measurements in the B-V domain hover around 20% depending on cometary activity, size, and composition (C. Opitom et al., 2025; A. D. Storrs et al., 1992; D. Jewitt & K. J. Meech, 1986).
This trend has two sources. First, reddening is defined as the reflectance spectral slope normalized by the mean reflectance, not the raw slope. This normalization causes wavelength-dependent variations even for linear reflectance spectra, with higher values at shorter wavelengths. Second, reflectance spectra are often non-linear and they steepen at shorter wavelengths, further increasing reddening values. For 3I/ATLAS, reflectance spectra in several studies also exhibit such steepening at short wavelengths (c.f. T. Santana-Ros et al., 2025; A. Alvarez-Candal et al., 2025).
To obtain reddening values at the V and UVW1 effective wavelengths for water production rate determination, we converted the middle panel’s reddening measurements from other wavelengths to UVOT wavelengths. For each measurement, we constructed a linear reflectance spectrum that, when multiplied by the solar spectrum, produces a red dust spectrum with the reported reddening value at that measurement wavelength. We then forward-modeled this reddened spectrum through Swift’s UVW1 and V filters and computed the reddening at UVOT effective wavelengths. Uncertainties in the converted reddening values were derived using error propagation.
We present the converted results in the right panel of Fig. 2. The y-axis still shows the original measurement wavelengths, while the x-axis displays the reddening converted to UVOT wavelengths. All reddening values increased after conversion, as expected from the reddening definition. Nearly all converted values exceed 20%, with a mean of 30.3%. At the 20% lower limit, the water production rate is molecules/s (SNR = 2.1), while at the mean reddening value, it reaches with SNR = 3.8, indicating clear detection. Our calculations assume linear reflectance spectra. When considering that reflectance spectra will steepen towards shorter wavelengths, actual reddening values would be higher, strengthening the water detection. The conversion accuracy depends on the wavelength difference between the original measurement (y-axis) and UVOT’s middle effective wavelength (4381.8 Å) — smaller differences yield higher accuracy. The measurement by A. Alvarez-Candal et al. (2025), nearest to our UVOT filter set (3200-5600 Å), converts to 38.6%, corresponding to a definitive detection of with SNR = 5.0, which corresponds to 40 kg s-1 at a heliocentric distance of 3.51 au.
The ratio log[Af/Q(H2O)] = -24.9 places 3I/ATLAS among the most dust-rich comets observed, comparable to a small subset that includes C/1995 O1 (Hale-Bopp), C/1991 B1 (Shoemaker-Levy), and C/2009 P1 (Garradd) (M. F. A’Hearn et al., 1995; T. Farnham et al., 1997; Y. R. Fernandez, 2000; D. Bodewits et al., 2014). We note, however, that the Af values in the M. F. A’Hearn et al. (1995) sample were not corrected for phase angle and OH production rates were used there, which will be lower than water production rates.
Using a sublimation model222https://pdssbn.astro.umd.edu/tools/ma-evap/index.shtml (J. J. Cowan & M. F. A’Hearn, 1979) and assuming constant solar elevation (as would occur for slow rotation or a Sun-facing pole), a Bond albedo of 0.05%, and unit infrared emissivity, we estimate the minimum required active area to be 19.47 km2, equivalent to a circular area with radius 2.47 km.
Observations acquired with the Hubble Space Telescope placed an upper limit on the nucleus’ radius of 2.8 km (D. Jewitt et al., 2025). If the water is produced directly by sublimation of ice within the nucleus, it would require that 20% of the surface contributes, whereas most solar system comets have active fractions around 3-5% (M. F. A’Hearn et al., 1995; H. U. Keller et al., 2015).
Gas detections beyond 2.5 au are uncommon due to the inefficiency of water sublimation at those distances. Only a few comets, such as C/1980 E1 (Bowell) and C/2009 P1 (Garradd), have shown OH detections at such large distances. OH emission from Bowell was detected at 5.3 au (M. F. Ahearn et al., 1984), requiring an implausibly large active area (1.6 107 km2) unless icy grains contributed significantly to water release. The OH production showed non-monotonic evolution, suggesting transient grain-driven activity. Similarly, Swift and SOHO/SWAN observations of comet Garradd revealed elevated water production beyond 3.4 au compared to narrow-slit spectroscopy, consistent with extended sources such as icy grains in the coma (M. R. Combi et al., 2013; D. Bodewits et al., 2014).
Near-infrared spectroscopic observations of 3I/ATLAS obtained with Gemini South/GMOS on July 5, 2025 UTC, and IRTF/SpeX on July 14, 2025 UTC, suggest the presence of large icy grains in the coma (B. Yang et al., 2025). If these grains are the primary source of the water and OH observed in the coma, they may explain why OH (A – X) emission was detected before any CN emission. Typically, CN is among the most distant fragment species detected in comets due to the high fluorescence efficiency of its B – X (0,0) violet system near 3883 Å, which also benefits from lower atmospheric extinction and higher detector sensitivity compared to the OH emission near 3085 Å (c.f. A. L. Cochran & W. D. Cochran, 1991). Large, dark icy grains in the coma are more efficiently heated than the nucleus, leading to preferential loss of their more volatile components. A similar process was observed at 67P/Churyumov–Gerasimenko, where the transport of material from the southern to the northern hemisphere resulted in enhanced H2O abundances relative to more volatile species such as CO2, CO, and HCN, implying that these more volatile ices were lost as lofted backfall material (H. U. Keller et al., 2017).
The physical processes that govern distant cometary activity and the retention or release of volatiles in such regimes remain poorly understood. Observational biases, limited statistics, and the challenge of detecting tenuous outgassing hinder progress. However, the unusual characteristics of 3I/ATLAS including early OH but no CN detection, strong coma reddening, and inferred large grains, highlight the need for long-term, multiwavelength monitoring. Simultaneous observations of gas and ice (e.g., via JWST) will be essential to disentangle the roles of native nucleus sublimation and extended grain activity, as well as probe the abundances of any hypervolatiles like CO and CO2. As noted by both M. J. Hopkins et al. (2025) and A. G. Taylor & D. Z. Seligman (2025), 3I/ATLAS may be water-rich, consistent with planetesimals formed in the outer regions of low-metallicity planetary systems. Our results offer one of the first direct tests of that hypothesis, finding indirect evidence for a large H2O production rate, but it remains to be seen if 3I/ALTAS’ coma is dominated by the other volatiles.
The presence of substantial OH, and the inferred production of H2O, can then be used in direct comparison to 2I/Borisov. Borisov had a much lower intercept speed with our solar system and likely came from a more recently formed system; the high CO/H2O ratio post-perihelion was seen as evidence of formation near the CO ice line, likely of an M-type star (D. Bodewits et al., 2020). 3I/ATLAS, which is moving nearly twice as fast as Borisov, is also likely to show asymmetric activity around perihelion as the thermal wave penetrates into deeper reservoirs. The post-perihelion behavior of 3I will expose more pristine materials, unaltered from the inbound journey. 3I/ATLAS has a H2O production rate at 3.5 au similar to the maximum value measured for 2I/Borisov at 2.4 au (J. Crovisier et al., 2019; Z. Xing et al., 2020), and therefore we propose the following tests of the H2O-rich hypothesis:
1.
If 3I/ATLAS is H2O-rich, indicating an older age and formation in a low-metallicity system, then the H2O production rate will peak near perihelion, and only trace amounts of high-metallicity volatiles like CO,CO2, CN and hydrocarbon species will be found.
2.
If 3I/ATLAS is hyper-volatile rich (CO/CO2), indicating that its dynamic age does not necessarily represent the metallicity of where it formed, the H2O production rate will drop precipitously postperihelion as more volatile species are sublimated from depth. This would be similar to the CO/H2O observations of 2I/Borisov (D. Bodewits et al., 2020).
If the first hypothesis is born out, and 3I/ATLAS is indeed from an older low-metallicity system, than this may suggest that there is a part of the planetesimal formation or removal process that may have ejected more mass from forming systems in order to properly place the large size of 3I into context (A. G. Taylor & D. Z. Seligman, 2025). Given the ISO production of low-metallicity systems would likely come from planetary interactions (S. N. Raymond et al., 2018) or stripping by stellar flybys (S. Pfalzner et al., 2021) an overabundance of large, H2O-rich ISO’s could serve as a key indicator of ejection conditions.
4 Conclusion
We report the detection of OH (A – X) emission in the coma of 3I/ATLAS using ultraviolet imaging with the Neil Gehrels-Swift Observatory. This confirms the presence of water in the third known interstellar object. The derived production rate of molecules s-1 (40 kg s-1) is notable given the object’s large heliocentric distance of 3.51 au, where water sublimation is typically inefficient. This level of activity is consistent with a reasonably-sized active area assuming equilibrium sublimation, but could also result from the presence of large icy grains in the coma, as suggested by contemporaneous near-infrared observations. If the coma of 3I/ATLAS remains H2O dominated and depleted in higher-metallicity volatiles like CO, CN, and CO2, the relatively large size of the interstellar object may be a telling clue about the planetesimal and mass loss processes of early low-metallicity systems.
The detection of OH emission prior to any detection of CN is unusual and may indicate differences in grain composition or volatile distribution compared to solar system comets (C. Opitom et al., 2025; A. Alvarez-Candal et al., 2025). While similar behavior has been observed in comets such as C/1980 E1 (Bowell) and C/2009 P1 (Garradd), the processes governing distant activity and the release of volatiles remain poorly constrained. Grain-driven activity, in particular, may play a more significant role in interstellar objects than previously assumed. This detection demonstrates the critical role of UV space-based observations in tracking the onset and evolution of activity in distant or dynamically unusual objects. Continued monitoring of 3I/ATLAS will help clarify the thermal and compositional properties of interstellar ices and improve our understanding of volatile retention and release in planetary systems beyond our own.